# Optimal Trudinger-Moser inequalities on complete noncompact Riemannian manifolds: Revisit of the argument from the local inequalities to global ones

### v2

@@ 3. Green’s function estimate
[…]
where $r=d(x,y)$ and $|\widetilde{g}|=\textit{det}(\widetilde{g}_{ij})$ denotes the determinant of the corresponding metric tensor. By carefully choosing the smooth cut-off function $f$, both the first two terms $|(n-3)f^{\prime}-rf^{\prime\prime}|$ and the coefficient of the third term $((n-2)f-rf^{\prime})$ can be uniformly bounded by $C(n,\delta,i)$. With the help of Lemma 3.1, we only need to consider the integral of the form
$\int_{B_{\delta}(x_{j})}|\nabla_{z}\ln\sqrt{|\widetilde{g}|}|(z)H(z,y)dV_{z}.$
An easy change of variablse from polar coordinates to harmonic coordinates shows
$\widetilde{g}_{ij}=g_{kl}\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\frac{\partial x^{l}}{\partial\widetilde{x}^{j}},$
where $\{x^{k}\}$ and $\{\widetilde{x}^{i}\}$ denotes the coordinate variables respectively. For simplicity we denote the transition matrix $A=\left(\frac{\partial x^{k}}{\partial\widetilde{x}^{i}}\right)$ then
$|\nabla\ln\sqrt{|\widetilde{g}|}|\leq\frac{|\nabla|g||}{2|g|}+\frac{|\nabla|A||}{|A|}.$
Since in harmonic coordinates $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$, we have
$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|g||}{|g|}(z)H(z,y)dV_{z}\right|$ $\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$
Similar to the argument in the proof of Lemma 3.1, the integral can be splitted into 3 pieces. Here we just check the part over $B_{\rho}(x)\cap B_{\delta}(x_{j})$, the other two parts can be treated in a similar manner:
$\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}\frac{|\nabla g(z)|}{d(z,y)^{n-2}}dV_{z}$ $\leq\frac{C(n)}{d(x,y)^{n-2}}\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}|\nabla g|dV_{z};$
[…]
$\leq\frac{C(n,Q)}{d(x,y)^{n-2-\frac{n}{p^{\prime}}}}.$
On the other hand, $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ implies there exists $C(n,Q)>0$ such that $|A|\geq C(n,Q)>0$. Hence we have
$\left|\int_{B_{\delta}(x_{j})}\frac{|\nabla|A||}{|A|}(z)H(z,y)dV_{z}\right|\leq C(n,Q)\int_{B_{\delta}(x_{j})}\frac{|\nabla|A|(z)|}{d(z,y)^{n-2}}dV_{z}$
We can estimate the above integral in a similar manner, but this time we make use of the $W^{2,p}$ estimate of the harmonic coordinate function $x^{i},i=1,\dots,n$. Indeed, in the derivation of the existence of harmonic coordinates, the $L^{p}$ theory of elliptic operators gives the uniform bound of $x^{i}||_{W^{2,p}}$, where $p$ satisfies $\alpha=1-\frac{n}{p}$. Hence there exists $C(n,\lambda,i,Q)$ such that $||\nabla A|_{L^{p}}\leq C(n,\lambda,i,Q)$.
Therefore from Lemma 3.1, $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a constant $C=C(n,\lambda,i,Q)$. To obtain the uniform estimate for Green’s function, it suffices to prove the uniform $C^{1}$ estimate of $F(x,y)$.
@@
Let $F(x,y)$ be a $W^{1,2}_{0}(B_{\delta}(x_{j}))$ solution of the Dirichlet problem (3.1), there exists a positive constant $C=C(n,\lambda,i,Q)$ such that $\nabla_{y}F(x,y)||_{\infty}\leq C$.
@@
4. Proof of Theorem 1.1
To prove Theorem 1.1, we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one. We first prove
@@
With the help of Green’s function estimates in the last section, the proof of Theorem 4.2 is exactly the same as that of Theorem 3.1 in [10] which is one of the main results. The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$. This shows the constants in the Moser-Trudinger inequality on any geodesic balls only depend on the radius $\delta$ which in turn depends on $\lambda$ and $i$.
Proof of Theorem 1.1.
For each $j$, we take cutoff function $\psi_{j}\in C_{0}^{\infty}(B_{\delta}(x_{j}))$ satisfying $0\leq\psi_{j}\leq 1$, $\psi_{j}=1$ on $B_{\frac{\delta}{2}}(x_{j})$ and $|\nabla\psi_{j}|\leq\frac{4}{\delta}$. Then $|\nabla\psi_{j}^{2}|=2\psi_{j}|\nabla\psi_{j}|\leq\frac{8}{\delta}\psi_{j}$ and $1\leq\sum_{j}\psi_{j}(x)\leq N$. Thus we have
(4.1) $\int_{M}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g}$
(4.2) $\leq$ $\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}(\alpha_{n}|u|^{\frac{n}{n-1}})dV_{g},$
(4.3) $\leq$ $\sum_{j}\int_{B_{\delta}(x_{j})}\phi_{n}(\alpha_{n}|\psi_{j}^{2}u|^{\frac{n}{n-1}})dV_{g},$
[…]
(4.6) $\leq$ $C(n,\lambda,i)2^{n}\sum_{j}\int_{M}\psi_{j}^{2n}|\nabla u|^{n}+|\nabla\psi_{j}^{2}|^{n}|u|^{n}dV_{g},$
(4.7) $\leq$ $C(n,\lambda,i)2^{n}\left(N\int_{M}|\nabla u|^{n}dV_{g}+\frac{8^{n}}{\delta^{n}}N\int_{M}|u|^{n}dV_{g}\right),$
(4.8) $\leq$ $C(n,\lambda,i)(N+N\frac{8^{n}}{\delta^{n}\tau^{n}}).$
This is a constant only depending on the initial geometry of the manifold.
Here we make some remarks about our alternative proof.
Remark 4.1.
To establish global inequalities from local ones, we choose appropriate cutoff functions $\psi_{j}$ such that $\nabla(\psi_{j}^{2}u)||_{L^{n}}\leq 1$. Thus the scale parameter $\delta$ actually depends on $\tau$. On the other hand, in the Green’s function estimates, we make use of both $W^{1,p}$ bounds of $g$ and $W^{2,p}$ bounds of $x^{i}$. There we need to choose proper value of $p$, hence $\alpha$. This can be done since the harmonic coordinates exists for arbitrary chosen $\delta$’s and $\alpha$’s. In other words, our inequalities can only be proved in specific harmonic coordinates.

### v3

@@ 2. Some preliminaries in differential geometry
$\Delta_{z}H(x,z)=\frac{1}{(n-2)\omega_{n-1}r^{n-1}}\left((n-3)f^{\prime}-rf^{\prime\prime}+((n-2)f-rf^{\prime})\left[\Delta_{z}d(x,z)-\frac{n-1}{d(x,z)}\right]\right),$
where $r=d(x,z)$. By carefully choosing the smooth cut-off function $f$, the first two terms $|(n-3)f^{\prime}-rf^{\prime\prime}|$ can be uniformly bounded by $C(n,\delta,i)$. With the help of Lemma 4.1, we need to consider the integral
$\left|\int_{B_{\delta}(x_{j})}\frac{1}{d(x,z)^{n-1}}((n-2)f-rf^{\prime})\left[\Delta_{z}d(x,z)-\frac{n-1}{d(x,z)}\right]\frac{1}{d(z,y)^{n-2}}dV_{z}\right|.$
First notice that $|(n-2)f-rf^{\prime}|\leq C(n,\delta,i)$. Applying Bishop comparison theorem and Günther comparison theorem,
$(n-1)a\cot ad(x,z)\leq\Delta_{d}(x,z)\leq(n-1)\sqrt{-\lambda}\coth\sqrt{-\lambda}d(x,z).$
Using Taylor expansions for $\cot t$ and $\coth t$, there exists a uniform constant $C=C(n,\delta,i,a,\lambda)$ such that $|\Delta_{z}d(x,z)-\frac{n-1}{d(x,z)}|\leq Cd(x,z)$. Therefore from Lemma 4.1, $\Gamma_{2}(x,y)$ is bounded by $d(x,y)^{n-3}$, up to some uniform constant. Then we argue inductively and hence conclude that $\Gamma_{k+1}$ together with its derivatives, are uniformly bounded by a uniform constant (for its derivatives, we leave the proof to readers). To obtain the uniform estimate for Green’s function, it suffices to prove the uniform $C^{1}$ estimate of $F(x,y)$.
@@
Let $F(x,y)$ be a $W^{1,2}_{0}(B_{\delta}(x_{j}))$ solution of the Dirichlet problem (4.1), there exists a positive constant $C=C(n,\lambda,a,i,Q)$ such that $\nabla_{y}F(x,y)||_{\infty}\leq C$.
@@
5. Alternative proof of Trudinger-Moser inequalities using the uniform estimates of Green functions
In this section, we give some remarks about how to prove Theorem 1.1 under additional condition of sectional curvature being upper bounded using the uniform Green function estimates. we achieve this through 2 steps: establishing the uniform Trudinger-Moser inequality on each $B_{\delta}(x_{j})$; gluing all inequalities from Step 1 so we extend our inequality from the local inequalities to a global one. We first prove
@@ 5. Alternative proof of Trudinger-Moser inequalities using the uniform estimates of Green functions
With the help of Green’s function estimates in the previous section, the proof of Theorem 5.2 is exactly the same as that of Theorem 3.1 in [11]. The parameters $\lambda,i$ do not explicitly appear in the estimates, but will determine $Q$ and $r_{H}(\alpha,0,Q)$, which gives the bounds of $\delta$. This shows the constants in the Moser-Trudinger inequality on any geodesic balls only depend on the radius $\delta$ which in turn depends on $\lambda$ and $i$.
Proof of Theorem 1.1 under sectional curvature upper bounded condition then follows from the uniform Green function estimates.